Consider an economy with two goods, cheese and salad. Suppose that the prices and quantities sold in 2012 and 2013 are the following:

Good

2012 Price

2013 Price

2012 Quantity

2013 Quantity

Cheese

10

12

10

9

Salad

15

16

5

8

a. Calculate the nominal GDP for 2012 and for 2013. Explain how you got your numbers.

b. Calculate the real GDP for 2013 using 2012 as a base year

c. Calculate the cost of the CPI basket if the market basket includes 2 cheese and 1 salad in 2012
Answer #1

a.

the nominal GDP for 2012

=(price*quantity)

=(10*10+15*5)

=175

the nominal GDP for 2013

=(price*quantity)

=(12*9+16*8)

=236

b.

Calculate the real GDP for 2013

=(price of 2012*quantity of 2013)

=(10*9+15*8)

=210

c.

the cost of the CPI basket in 2012=2*10+1*15=35

the cost of the CPI basket in 2013=2*12+1*16=40